# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

### v1

@@ 6.3. Pipeline for Creating Spatialized Vibrotactile Systems
Synthesizing the processes followed during the three case studies, herein we summarize a generalized five-step pipeline for creating spatialized vibrotactile systems.
@@
The first step is to determine what information to be delivered through vibrotactile feedback. For example, the purpose of the phonemic display was to convey vowels and consonants to deaf and hard-of-hearing people, so the types of vowels and consonants in words were the target information. In the drone teleoperation scenario, the key information was the potential risk of colliding with nearby obstacles and could be further divided into the directions of obstacles and their risk levels.
@@
The second step is to decide how to convey the information. Since vibrations can be configured with different characteristics such as intensity, frequency, duration, position, and so on, the common practice is to breakdown the information and deliver it through multiple channels to maximize information transmission (Tan et al., 2020). For example, vowels and consonants for the phonemic displays were distinguished by using moving patterns and static patterns, while different consonants were distinguished by actuator position, frequency modulation and vibration duration. As shown in prior research, providing information through multiple channels with fewer levels in each channel can enhance the overall information transfer for vibrotactile systems.
@@
The third step is to decide where to install the system and derive the layout of vibration units. In cases where the information was spatially connected to the surrounding environment, the layout should have a good mapping between tactile positions and spatial direction. For example, in the VR fitness application, vibration units were mainly placed on the arm to provide localized feedback, indicating the arm pose differences between the user and the wall. In the UAV teleoperation, locations of vibrations were perceptually mapped to the directions of incoming obstacle. Perceptual limits should also be considered, e.g., spatial acuity indicated the upper limits of layouts, as adding more actuators to the layout would not increase the information transmission.
@@
The fourth step is to build the system. Control units and vibration units can be chosen based on the scale of the system. Learning from the building process of case studies, We recommend using the small control unit for localized systems such as sleeves and headbands, as well as applications that require short-term usage. Alternatively, the large control unit is suitable for long-term usage and body-scale systems, such as immersive virtual experience and robot teleoperation. When designing vibration patterns, if the information involves complex waveforms, designers can utilize the waveform library in the GUI Editor to expedite the authoring process, just like how we designed patterns for the phonemic display. If the information only requires changing levels of intensity or frequency, the Python server with APIs should be handy.
@@
The final step is to test the system and iterate on the design. If the desired information cannot be transmitted as expected, the designer can reconfigure the system by changing the layout or experiment different characteristics for conveying the information.
6.4. Potential Applications for VibraForge
A toolkit such as VibraForge could benefit several domains. In VR, spatial vibrotactile feedback can enhance immersive social experiences. In VR social platforms, adding tactile interactions can help users express their emotions through social touch gestures (Sykownik and Masuch, 2020). If developers are creating devices for users with sensory issues in certain areas of their body, they can leverage the toolkit to quickly and easily adjust the layout of actuators to avoid discomfort.
During human-robot interaction, spatial feedback could be useful for assisting humanoid robot teleoperation. Humanoid robot training collects data from human operators performing tasks remotely. However, the absence of tactile feedback mechanisms makes it difficult for operators to execute daily tasks that are typically performed with ease using their own bodies (Pacchierotti and Prattichizzo, 2024). Spatial tactile feedback can provide additional information about remote interaction, such as when a robot hits a wall or other obstacle. The choice of tactile feedback positions can also be configured to match the positions of sensors on the robot. This feedback could thus increase an operator’s situational awareness and close the interaction loop.
Last but not least, we envision that VibraForge can be used to build custom prototypes for tactile perceptual studies, such as those measuring spatial acuity, information transfer, or other perceptual characteristics. For example, Park et al. performed perceptual studies to measure information transfer across five body positions (Park et al., 2023). These positions were not chosen by the researchers but instead were constrained by the design of the commercial system they used (i.e., tactile modules from bHaptics (bHaptics, 2024)). By using VibraForge, researchers could freely adjust the layout of vibration units on the body and adapt it to any test they want to conduct.
@@ 7. Conclusion
Acknowledging the limits with existing spatialized vibrotactile feedback systems and toolkits, we built the VibraForge toolkit. Our toolkit enables scalability via a modular design and the use of chain-connection data transmission methods. Technical evaluations validated the toolkit design, and case studies demonstrated the wide range of experiences that could be authored. The lessons learned during the case studies contributed to our understanding of a generalized pipeline for creating spatialized vibrotactile systems, and highlighted potential research directions of toolkit refinement and future applications. The toolkit is in its final stages of preparation, and will be released to the public later this year. We are excited to see how researchers and designers will harness the toolkit to elevate their own research and push the boundary of design innovation when using spatialized vibrotactile feedback.

### v2

@@ 7.3. Pipeline for Creating Spatialized Vibrotactile Systems
Synthesizing the processes followed during the case studies, we summarize a generalized five-step pipeline for creating spatialized vibrotactile systems. The first step is to determine what information to deliver using vibrotactile feedback. For example, the purpose of the phonemic display was to convey vowels and consonants to deaf and hard-of-hearing people, so the types of vowels and consonants in words were the target information.
@@
The second step is to decide how to convey the information. Since vibrations can be configured with different characteristics such as intensity, frequency, duration, position, and so on, the common practice is to break down the information and deliver it via multiple channels to maximize information transmission (Tan et al., 2020). For example, vowels and consonants for the phonemic displays were distinguished by using moving patterns and static patterns, while different consonants were distinguished by actuator position, frequency modulation and vibration duration.
@@
The third step is to decide where to install the system and derive the layout of vibration units. In cases where the information was spatially connected to the surrounding environment, the layout should have a good mapping between tactile positions and spatial direction. In the UAV teleoperation, locations of vibrations were perceptually mapped to the directions of incoming obstacle. Perceptual limits should also be considered, e.g., spatial acuity indicated the upper limits of layouts, as adding more actuators to the layout would not increase the information transmission.
@@
The fourth step is to build the system. Control and vibration units should be chosen based on the scale of the system (e.g., the small control unit for localized systems such as sleeves and headbands and the large control unit for long-term usage and body-scale systems such as immersive virtual experience and robot teleoperation). When designing vibration patterns, if the information involves complex waveforms, designers can utilize the Waveform Library in the GUI Editor to expedite the authoring process, similar to its use for the phonemic display. If the information only requires parameterized control, the Python server and APIs should be handy.
@@
The final step is to test the system and iterate on the design. If the desired information is not transmitted as expected, designers can consider changing the layout or using different vibration characteristics to convey the information.
@@ 8. Conclusion
The VibraForge toolkit enables scalability and expressivity via a modular design and a chain-connection method. Technical evaluation validated toolkit’s design requirements, case studies demonstrated the design space, and usability study highlighted low technical barrier and customizability. The lessons learned from evaluations contributed to potential directions for toolkit refinement and generalized pipeline for creating spatialized vibrotactile systems. We are excited to see how researchers and designers will harness the toolkit to elevate their own research and push the boundaries of design innovation using spatialized vibrotactile feedback.
@@ Acknowledgements.
We would like to express our deepest gratitude to Dr. Michelle Annett, whose assistance significantly enhanced the quality of our manuscript. We would also like to thank Jeb Thomas and Brenna Li for providing feedback on paper writing. We also thank Mengfei Liu, who kindly provided the sewing machine for tailoring the garments used in this work. Some figures in the manuscript were designed and improved by For Real Entertainment and Annie Squarecircle. ChatGPT was used in the writing process for language enhancement.